# The continuum limit of some products of random matrices associated with renewing flows

### v1

@@ 1.6. Outline of the paper and main results
(1.35) $\frac{1}{\tau^{2}}L^{\ast}(k,\ell)=\frac{3\left(\ell^{\prime}-2\right)^{2}}{8}+\frac{9\left(\ell^{\prime}-2\right)\left(\ell^{\prime}+2\right)}{20}k^{2}\\ -\frac{243(\ell^{\prime})^{4}-36504(\ell^{\prime})^{2}+21168}{56000}k^{4}\\ +\frac{59049(\ell^{\prime})^{6}-39724668(\ell^{\prime})^{4}+1597795632(\ell^{\prime})^{2}-421920576}{1601600000}k^{6}+\cdots$
Plots of $L$ and $L^{\ast}$ are shown in Figure 1 (b).
In §6, we show that, for $0\leq k^{2}\leq 1/d$, the function $\ell\mapsto L(k,\ell)$ is of intrinsic interest as the generalised Lyapunov exponent of (the continuum limit of) a product of random matrices associated with a renewing flow. Roughly speaking, this renewing flow is obtained from the original one by inserting time intervals of pure strain in the $x_{i}-x_{j}$ planes, $i<j$; the number $1/d-k^{2}$ may then be interpreted as a measure of the strength of the disorder associated with the pure-straining, relative to the disorder in the original velocity gradient.
@@ 2.1. The Lie algebra ${\mathfrak{s}l}(d)$
2.2. Proof of Proposition 1.1
We begin by observing that, by definition, the functions in $V_{\ell}$ are homogeneous of degree $\ell$, so that, for every $v\in V_{\ell}$, every $x\in{\mathbb{R}}_{\ast}^{d}$ and every $r>0$, there holds the identity
@@
2.3. Quasi-solvability
Here “quasi-solvability” refers to the well-known fact that there are special values of $\ell$ for which some part of the spectrum of the transfer operator may be obtained by purely algebraic means [15]. The explanation is that the infinitesimal generators map polynomials to polynomials. Hence the
@@ 2.3. Quasi-solvability
Let $\ell$ be an even natural number. Then the subspace $\mathcal{P}$$_{\ell}$ of $V_{\ell}$, consisting of the homogeneous polynomials of degree $\ell$, is invariant under the representation $T_{\ell}$, and hence also under the transfer operator $\mathscr{T}$$_{\ell}$ defined by Equation (1.16).
@@
This remains true in the continuum limit, and we shall find upon calculation that, when $\ell$ is an even natural number, there exist non-trivial subspaces of $\mathcal{P}$$_{\ell}$, of much lower dimension, that contain $1_{\ell}$ and are invariant under the operator (1.23).
@@ 3.1. The adjoint of the transfer operator
defined on $V_{-\overline{\ell}-d}$ is the formal adjoint of the transfer operator $\mathscr{T}$$_{\ell}$ defined on $V_{\ell}$ in the previous paragraph. Now, for our particular product, the random matrix $g$ has a very particular structure, displayed in Equation (1.6): its inverse is obtained by changing the sign of $\delta t$ and reversing the order in which the factors occur. In the continuum limit where we neglect terms of order $o(\tau^{2})$, neither the sign of $\delta t$ nor the order matters. Since the spectrum of the adjoint is the complex conjugate of the spectrum, we deduce the
@@
$\lambda$ is an eigenvalue of the operator $\mathscr{T}$$_{\ell}$, defined by Equation (1.23), if and only if its complex conjugate is an eigenvalue of ${\mathscr{T}}_{-\overline{\ell}-d}$. In particular, if $\ell$ and $\lambda(\ell)$ are real, then
(3.6) $L(\ell)=L(-\ell-d)\;\;\text{and}\;\;L^{\ast}(\ell)=L^{\ast}(-\ell)-\ell\,d\,.$
@@ 4.2. Expansion in powers of $k^{2}$
(4.27) $\mu=\mu_{0}:=-\frac{\ell\left(\ell+2\right)}{2}k^{2}+\frac{\left(\ell-2\right)\ell\left(\ell+2\right)\left(\ell+4\right)}{128}k^{4}+\frac{\left(\ell-2\right)\ell\left(\ell+2\right)\left(\ell+4\right)}{256}k^{6}\\ -\frac{\left(\ell-2\right)\ell\left(\ell+2\right)\left(\ell+4\right)\left(7\ell^{4}+28\ell^{3}+36\ell^{2}+16\ell-5248\right)}{2097152}k^{8}+\cdots$
This leads to the expansions (1.32-1.33). By re-arranging the terms in the former series and collecting them in powers of $\ell$, we deduce expansions for all the cumulants:
(4.28) $\frac{1}{\tau^{2}}\,\gamma_{1}=1-k^{2}-\frac{1}{8}k^{4}-\frac{1}{16}k^{6}-\frac{41}{1024}k^{8}-\frac{59}{2048}k^{10}+\cdots$
(4.29) $\frac{1}{\tau^{2}}\,\frac{\gamma_{2}}{2}=\frac{1}{2}-\frac{1}{2}k^{2}-\frac{1}{32}k^{4}-\frac{1}{64}k^{6}-\frac{81}{8192}k^{8}-\frac{115}{16384}k^{10}+\cdots$
(4.30) $\frac{1}{\tau^{2}}\,\frac{\gamma_{3}}{6}=\frac{1}{32}k^{4}+\frac{1}{64}k^{6}+\frac{169}{16384}k^{8}+\frac{251}{32768}k^{10}+\cdots$
(4.31) $\frac{1}{\tau^{2}}\,\frac{\gamma_{4}}{24}=\frac{1}{128}k^{4}+\frac{1}{256}k^{6}+\frac{361}{131072}k^{8}+\frac{571}{262144}k^{10}+\cdots$
@@ 4.3. Quasi-solvability
The reader will easily verify that these small-$k^{2}$ expansions do agree with the formulae presented in the previous paragraph. These exact results also make clear that the radius of convergence decreases with $\ell$; in particular, for $\ell\geq 6$, we expect that the expansions do not converge for $k^{2}=1/2$.
@@ 5.1. The small-$k^{2}$ limit
(5.10) $\mu=\mu_{0}:=-\frac{4\ell\left(\ell+3\right)}{5}k^{2}+\frac{12\ell\left(\ell-2\right)\left(\ell+3\right)\left(\ell+5\right)}{875}k^{4}\\ -\frac{72\left(\ell-2\right)\ell\left(\ell+3\right)\left(\ell+5\right)\left(9\ell^{2}+27\ell-395\right)}{3128125}+\cdots$
@@
(5.11) $\frac{1}{\tau^{2}}\,\gamma_{1}=2-\frac{12}{5}k^{2}-\frac{72}{175}k^{4}-\frac{34128}{125125}k^{6}\\ -\frac{17244576}{74449375}k^{8}-\frac{223736256}{1010384375}k^{10}+\cdots$
(5.12) $\frac{1}{\tau^{2}}\,\frac{\gamma_{2}}{2}=\frac{2}{3}-\frac{4}{5}k^{2}-\frac{12}{875}k^{4}+\frac{5976}{625625}k^{6}\\ +\frac{59358528}{2605728125}k^{8}+\frac{22326336}{642971875}k^{10}+\cdots$
(5.13) $\frac{1}{\tau^{2}}\,\frac{\gamma_{3}}{6}=\frac{72}{875}k^{4}+\frac{27432}{446875}k^{6}+\frac{30362256}{521145625}k^{8}+\frac{15082522656}{247544171875}k^{10}+\cdots$
(5.14) $\frac{1}{\tau^{2}}\,\frac{\gamma_{4}}{24}=\frac{12}{875}k^{4}+\frac{1584}{284375}k^{6}+\frac{32383188}{13028640625}k^{8}+\frac{60775592}{1237720859375}k^{10}+\cdots$
@@ 6. A renewing flow with intervals of pure strain
In §1.6, we introduced the artificial parameter $k^{2}$ in order to take advantage of Proposition 1.1, which relates the transfer operator to the Casimir of $\text{\rm SO}(d,{\mathbb{R}})$ when $k^{2}=1/d$. Our purpose in the present section is to show that, for every $0\leq k^{2}\leq 1/d$, $\ell\mapsto L(k,\ell)$ may itself be interpreted as the generalised Lyapunov exponent of (the continuum limit of) a product of random matrices associated with a renewing flow.
@@
where the $\xi_{ij}^{n}$ are given by the formulae (1.7-1.9) and the $\alpha_{ij}^{n}$ are mutually independent, zero-mean random variables independent of the $\xi_{ij}^{n}$. In the continuum limit, the transfer operator associated with the product becomes
@@
(a)(b)
Figure 2. Plots of the cumulants $\gamma_{j}$ against $1/d-k^{2}$ for $\tau=1$ and $1\leq j\leq 4$. The blue and red curves correspond to $d=2$ and $d=3$ respectively: (a) $\gamma_{1}$ (solid curves) and $\gamma_{2}$ (dashed curves); (b) $\gamma_{3}$ (solid curves) and $\gamma_{4}$ (dashed curves).

### v2

@@ 1.8. Outline of the paper and main results
(1.43) $\Upsilon(k,s)=\frac{3\left(s-2\right)^{2}}{8}+\frac{9\left(s-2\right)\left(s+2\right)}{20}k^{2}\\ -\frac{243s^{4}-36504s^{2}+21168}{56000}k^{4}\\ +\frac{59049s^{6}-39724668s^{4}+1597795632s^{2}-421920576}{1601600000}k^{6}+\cdots$
Plots of $\Lambda(1/\sqrt{3},\ell)$ against $\ell$ and $\Upsilon(1/\sqrt{3},s)$ against $s$ are shown in Figure 1 (b).
In §6, we show that, for $0\leq k^{2}\leq 1/d$, the function $\ell\mapsto\tau^{2}\Lambda(k,\ell)$ is of intrinsic interest as the generalised Lyapunov exponent of (the continuum limit of) a product of random matrices associated with a renewing flow. Roughly speaking, this renewing flow is obtained from the original one by inserting time intervals of pure strain in the $x_{i}-x_{j}$ planes, $i<j$; the number $1/d-k^{2}$ may then be interpreted as a measure of the strength of the disorder associated with the pure-straining, relative to the disorder in the original velocity gradient.
@@ 2.1. The Lie algebra ${\mathfrak{s}l}(d)$
2.2. The operator $\sum_{i\neq j}N_{ij}^{2}$ as a perturbation of the Casimir of $SO(d,{\mathbb{R}})$
We are now ready to carry out the proof of Proposition 1.1. We begin by observing that, by definition, the functions in $V_{\ell}$ are homogeneous of degree $\ell$, so that, for every $v\in V_{\ell}$, every $x\in{\mathbb{R}}_{\ast}^{d}$ and every $r>0$, there holds the identity
@@ 2.2. The operator $\sum_{i\neq j}N_{ij}^{2}$ as a perturbation of the Casimir of $SO(d,{\mathbb{R}})$
2.3. The replica trick
This “trick” refers to the well-known fact that there are special values of $\ell$ for which some part of the spectrum of the transfer operator may be obtained by purely algebraic means; see for instance [28] or [8], §2.4.2. The explanation is that the infinitesimal generators map polynomials to polynomials. Hence the
@@ 2.3. The replica trick
Let $\ell$ be an even natural number. Then the subspace $\mathcal{P}$$_{\ell}$ of $V_{\ell}$, consisting of the homogeneous polynomials of degree $\ell$, is invariant under the representation $T_{\ell}$, and hence also under the transfer operator $\mathscr{T}$$_{\ell}$ defined by Equation (1.23) and the operator $\Delta_{K}-k^{2}\sum_{i<j}A_{ij}^{2}$ in Equation (1.32).
@@
This remains true in the continuum limit leading to the spectral problem (1.32), where the object is to compute $\mu(k,\ell)$, and we shall find upon inspection that, when $\ell$ is an even natural number, there exist non-trivial subspaces of $\mathcal{P}$$_{\ell}$, of much lower dimension, that contain $1_{\ell}$ and are invariant under the operator $\Delta_{K}-k^{2}\sum_{i<j}A_{ij}^{2}$.
@@ 3.1. The fluctuation theorem
defined on $V_{-\overline{\ell}-d}$ is the formal adjoint of the transfer operator $\mathscr{T}$$_{\ell}$ defined on $V_{\ell}$ in the previous paragraph.
Now, for our particular product, the random matrix $g$ has a very particular structure, displayed in Equation (1.13): its inverse is obtained by changing the sign of $\delta t$ and reversing the order in which the factors occur. In the continuum limit where we neglect terms of order $o(\tau^{2})$, neither the sign of $\delta t$ nor the order matters. Since the spectrum of the adjoint is the complex conjugate of the spectrum, we deduce the
@@
For $k=1/\sqrt{d}$, there holds
(3.6) $\Lambda\left(k,\ell\right)=\Lambda\left(k,-\ell-d\right)\;\;\text{and}\;\;\Upsilon(k,s)=\Upsilon(k,-s)-s\,d\,.$
The proposition and its proof are essentially taken from Vanneste’s Proposition 2 [28]. It is evident from the formulae presented in §1.8 that the symmetry property (3.6) actually holds for every $k$. It is a particular instance of the fluctuation theorem which is well-known in non-equilibrium statistical mechanics; see [9], §10 and [26], §11.3.3.
@@ 4.1. Calculation of $\Lambda(k,\ell)$ to $o(\ell^{2})$
Remark 4.1.
Elliptic functions are widely used in analytic number theory, and this leads to interesting alternative ways of expressing this last formula. For instance, if we introduce the divisor function (see [22], §27.2)
$\sigma_{\alpha}(j):=\sum_{n|j}n^{\alpha}$
where the sum is over the divisors of the natural number $j$, then it may be shown that
$\sum_{j=1}^{\infty}\left(\frac{q^{2j}}{1-q^{4j}}\right)^{2}=\sum_{j=1}^{\infty}\sigma_{1}(j)\,q^{4j}$
and
$\sum_{j=1}^{\infty}\left(\frac{q^{2j-1}}{1-q^{4j-2}}\right)^{2}=\sum_{j=1}^{\infty}\sigma_{1}(j)\,q^{2j}-\sum_{j=1}^{\infty}\sigma_{1}(j)\,q^{4j}\,.$
The sum of the power series appearing on the right-hand side may be expressed in terms of the Eisenstein series $G_{2}$; see Apostol [1], §3.10.
@@ 4.2. Expansion in powers of $k^{2}$
(4.27) $\mu=\mu_{0}:=-\frac{\ell\left(\ell+2\right)}{2}k^{2}\\ +\frac{\left(\ell-2\right)\ell\left(\ell+2\right)\left(\ell+4\right)}{128}k^{4}+\frac{\left(\ell-2\right)\ell\left(\ell+2\right)\left(\ell+4\right)}{256}k^{6}\\ -\frac{\left(\ell-2\right)\ell\left(\ell+2\right)\left(\ell+4\right)\left(7\ell^{4}+28\ell^{3}+36\ell^{2}+16\ell-5248\right)}{2097152}k^{8}+\cdots$
This leads to the expansions (1.40-1.41). By re-arranging the terms in the former series and collecting them in powers of $\ell$, we deduce expansions for all the cumulants:
(4.28) $\gamma_{1}=1-k^{2}-\frac{1}{8}k^{4}-\frac{1}{16}k^{6}-\frac{41}{1024}k^{8}-\frac{59}{2048}k^{10}+\cdots$
(4.29) $\frac{\gamma_{2}}{2}=\frac{1}{2}-\frac{1}{2}k^{2}-\frac{1}{32}k^{4}-\frac{1}{64}k^{6}-\frac{81}{8192}k^{8}-\frac{115}{16384}k^{10}+\cdots$
(4.30) $\frac{\gamma_{3}}{6}=\frac{1}{32}k^{4}+\frac{1}{64}k^{6}+\frac{169}{16384}k^{8}+\frac{251}{32768}k^{10}+\cdots$
(4.31) $\frac{\gamma_{4}}{24}=\frac{1}{128}k^{4}+\frac{1}{256}k^{6}+\frac{361}{131072}k^{8}+\frac{571}{262144}k^{10}+\cdots$
@@ 4.3. The replica trick
The reader will easily verify that these small-$k^{2}$ expansions do agree with the formulae presented in the previous paragraph. These exact results indicate that the radius of convergence decreases with $\ell$; in particular, for $\ell\geq 6$, we expect that the expansions do not converge for $k^{2}=1/2$. The singularities on the circle of convergence occur at complex values of $k$ such that the “spectral gap” separating the dominant eigenvalue branch from the others shrinks to zero.
4.4. The band-center anomaly for Anderson’s model at zero energy
For $k^{2}=1/2$, Formula (1.36) yields
(4.37) $\gamma_{1}\left(\frac{1}{\sqrt{2}}\right)=2\frac{\boldsymbol{\mathsf{E}}\left(\frac{1}{\sqrt{2}}\right)}{\boldsymbol{\mathsf{K}}\left(\frac{1}{\sqrt{2}}\right)}-1=4\left[\frac{\Gamma(\frac{3}{4})}{\Gamma(\frac{1}{4})}\right]^{2}=\frac{1}{\pi G^{2}}$
where $G$ denotes the Gauss lemniscate constant. The same (apart from a trivial factor) Lyapunov exponent comes out of the analysis of a one-dimensional Anderson model which exhibits the Kappus–Wegner band-center anomaly at zero energy; see [10, 17, 24]. The formula also appears explicitly in [4], §6.2, and [23] in connection with products whose terms are matrices of the particular form
(4.38) $\text{SL}(2,{\mathbb{R}})\ni g_{n}=e^{\alpha_{n}K_{12}}\,e^{w_{n}A_{12}}$
where $\alpha_{n}$ and $w_{n}$ are independent random variables with zero mean and respective variances $D_{\alpha\alpha}$ and $D_{ww}$. More precisely, Comtet et al. found that, in the limit as both variances tend to zero, the Lyapunov exponent of the product equals
(4.39) $\Omega:=\left(D_{\alpha\alpha}+D_{ww}\right)\frac{\boldsymbol{\mathsf{E}}\left(k\right)}{\boldsymbol{\mathsf{K}}\left(k\right)}-D_{\alpha\alpha}$
where
(4.40) $k^{2}:=\frac{D_{ww}}{D_{\alpha\alpha}+D_{ww}}\,.$
In their study of the one-dimensional Dirac equation with a random mass, Ramola & Texier showed how to map the Anderson model at zero energy to a product of precisely this form; see [23], §8.2. There only remains to clarify the relationship between the products studied in the present paper and those considered by Comtet et al. in their §6.2.
To this end, let $k$ be defined by Formula (4.40), so that
(4.41) $\frac{k}{k^{\prime}}:=\frac{k}{\sqrt{1-k^{2}}}=\sqrt{\frac{D_{ww}}{D_{\alpha\alpha}}}\,.$
Denote by $\widetilde{L}(k,\ell)$ the generalised Lyapunov exponent associated with Comtet et al.’s product. In the continuum limit, it is given by the formula
$\widetilde{L}(k,\ell)=\widetilde{\tau}^{2}\,\widetilde{\Lambda}(k,\ell)+o\left(\widetilde{\tau}^{2}\right)\;\;\text{as $\widetilde{\tau}\rightarrow 0+$}$
where $\widetilde{\tau}^{2}=D_{\alpha\alpha}/2$ and $\widetilde{\Lambda}(k,\ell)$ is the leading eigenvalue of the operator
$K_{12}^{2}+\frac{D_{ww}}{D_{\alpha\alpha}}A_{12}^{2}\,.$
Recalling Equations (1.32) and (4.41), we deduce the relationship
$\widetilde{\Lambda}(k,\ell)=\mu\left(\text{\rm i}\frac{k}{k^{\prime}},\ell\right)$
where $\mu(\cdot,\ell)$ was calculated in the foregoing subsections. By setting $n=0$ in Equation (4.12), it follows in particular that
(4.42) $\widetilde{\tau}^{2}\frac{\partial\widetilde{\Lambda}}{\partial\ell}(k,0)=\frac{D_{\alpha\alpha}}{2}\,2\left[\frac{\boldsymbol{\mathsf{E}}\left(\text{\rm i}\frac{k}{k^{\prime}}\right)}{\boldsymbol{\mathsf{K}}\left(\text{\rm i}\frac{k}{k^{\prime}}\right)}-1\right]=D_{\alpha\alpha}\left[\frac{1}{(k^{\prime})^{2}}\frac{\boldsymbol{\mathsf{E}}\left(k\right)}{\boldsymbol{\mathsf{K}}\left(k\right)}-1\right]\\ =D_{\alpha\alpha}\left[\frac{D_{\alpha\alpha}+D_{ww}}{D_{\alpha\alpha}}\frac{\boldsymbol{\mathsf{E}}\left(k\right)}{\boldsymbol{\mathsf{K}}\left(k\right)}-1\right]=\Omega$
where, to obtain the second equality, we have made use of Formula 19.7.2 in [22] :
$\boldsymbol{\mathsf{E}}\left(\text{\rm i}\frac{k}{k^{\prime}}\right)=\frac{1}{k^{\prime}}\boldsymbol{\mathsf{E}}\left(k\right)\;\;\text{and}\;\;\boldsymbol{\mathsf{K}}\left(\text{\rm i}\frac{k}{k^{\prime}}\right)=k^{\prime}\boldsymbol{\mathsf{K}}\left(k\right)\,.$
Our calculations therefore translate into results pertaining to products involving matrices of the form (4.38). It is interesting to note that the case of equal variances, relevant to the Kappus–Wegner band-center anomaly, lies precisely on the circle of convergence of the expansion presented in §4.2.
4.5. Relationship to the Kraichnan model of §1.2
As we have already indicated, formulae like Equations (1.36-1.37) were found by Chetrite et al. for the Kraichnan model. Indeed, the analysis in their §4.5 leads to the following conclusion: the rate function associated with the two-dimensional Kraichnan model may be expressed, in their notation, as the Legendre transform of
(4.43) $\nu\mapsto(2\alpha+\beta+\gamma)\nu(\nu+1)-2(\alpha+\gamma)E_{\nu,0}$
where $E_{\nu,0}$ is the dominant eigenvalue of the operator
(4.44) ${\mathcal{L}}_{\nu}:=2\alpha\left(\sin\phi\frac{\partial}{\partial\phi}-\nu\cos\phi\right)^{2}+2\gamma\frac{\partial^{2}}{\partial\phi^{2}}$
acting on $\pi$-periodic functions. The substitution $\phi=2\theta$, $\nu=\ell/2$ brings this operator into the form (see Table 1)
$\frac{\gamma}{2}K_{12}^{2}+\frac{\alpha}{2}A_{12}^{2}\,.$
In their calculation, both $\alpha$ and $\gamma$ assume positive values. The Kraichnan model therefore maps directly to the product studied by [4], §6.2, and hence to ours. By the calculation of the previous paragraph, we deduce
$E_{\nu,0}:=\frac{\gamma}{2}\,\mu\left(\text{\rm i}\sqrt{\frac{\alpha}{\gamma}},\,2\nu\right)\,.$
We remark that, in the incompressible case $\beta=0$, Formula (1.8) for the degree of anisotropy of the Kraichnan model yields
$\kappa=\frac{\alpha}{\gamma}\,.$
Hence the case of “maximal anisotropy” lies on the circle of convergence of the expansion presented in §4.2.
@@ 5.1. The small-$k^{2}$ limit
(5.10) $\mu=\mu_{0}:=-\frac{4\ell\left(\ell+3\right)}{5}k^{2}+\frac{12\ell\left(\ell-2\right)\left(\ell+3\right)\left(\ell+5\right)}{875}k^{4}\\ -\frac{72\left(\ell-2\right)\ell\left(\ell+3\right)\left(\ell+5\right)\left(9\ell^{2}+27\ell-395\right)}{3128125}k^{6}+\cdots$
@@
(5.11) $\gamma_{1}=2-\frac{12}{5}k^{2}-\frac{72}{175}k^{4}-\frac{34128}{125125}k^{6}\\ -\frac{17244576}{74449375}k^{8}-\frac{223736256}{1010384375}k^{10}+\cdots$
(5.12) $\frac{\gamma_{2}}{2}=\frac{2}{3}-\frac{4}{5}k^{2}-\frac{12}{875}k^{4}+\frac{5976}{625625}k^{6}\\ +\frac{59358528}{2605728125}k^{8}+\frac{22326336}{642971875}k^{10}+\cdots$
(5.13) $\frac{\gamma_{3}}{6}=\frac{72}{875}k^{4}+\frac{27432}{446875}k^{6}+\frac{30362256}{521145625}k^{8}+\frac{15082522656}{247544171875}k^{10}+\cdots$
(5.14) $\frac{\gamma_{4}}{24}=\frac{12}{875}k^{4}+\frac{1584}{284375}k^{6}+\frac{32383188}{13028640625}k^{8}+\frac{60775592}{1237720859375}k^{10}+\cdots$
@@ 6. A renewing flow with intervals of pure strain
In §1.8, we introduced the artificial parameter $k^{2}$ in order to take advantage of Proposition 1.1, which relates the transfer operator to the Casimir of $\text{\rm SO}(d,{\mathbb{R}})$ when $k^{2}=1/d$. Our purpose in the present section is to show that, for every $0\leq k^{2}\leq 1/d$, $\ell\mapsto\Lambda(k,\ell)$ may itself be interpreted as the generalised Lyapunov exponent of (the continuum limit of) a product of random matrices associated with a renewing flow.
@@
where the $\xi_{ij}^{n}$ are given by the formulae (1.14-1.16) and the $\alpha_{ij}^{n}$ are mutually independent, zero-mean random variables independent of the $\xi_{ij}^{n}$. In the continuum limit, the transfer operator associated with the product becomes
@@
Refer to caption(a)(b)
Figure 2. Plots of the cumulants $\gamma_{j}$ against $1/d-k^{2}$ for $1\leq j\leq 4$. The blue and red curves correspond to $d=2$ and $d=3$ respectively: (a) $\gamma_{1}$ (solid curves) and $\gamma_{2}$ (dashed curves); (b) $\gamma_{3}$ (solid curves) and $\gamma_{4}$ (dashed curves).
The fact that, in the case $d=2$, the first two cumulants nearly coincide is reminiscent of the “single parameter scaling” property exhibited by other models; see [25].